I need to determine if this proportion is true or false:
3/6 = 6/x

Answer by tutorcecilia(2152)

3/6 = 6/x
(3)(x) / (6)(6) [ Cross-multiply]
3x = 36 [Solve for "x"]
3x/3 = 36/3
x = 12
So, the equation is true when x=12.